# Growth and collapse of subsystem complexity under random unitary circuits

### v1

@@ Abstract
For chaotic quantum dynamics modeled by random unitary circuits, we study the complexity of reduced density matrices of subsystems as a function of evolution time where the initial global state is a product state. The state complexity is defined as the minimum number of local quantum channels to generate a given state from a product state to a good approximation. In 1+11+1d, we prove that the complexity of subsystems of length $\ell$ smaller than half grows linearly in time $T$ at least up to $T=\ell/4$ but becomes zero after time $T=\ell/2$ in the limit of a large local dimension, while the complexity of the complementary subsystem of length larger than half grows linearly in time up to exponentially late times. Using holographic correspondence, we give some evidence that the state complexity of the smaller subsystem should actually grow linearly up to time $T=\ell/2$ and then abruptly decay to zero.
@@ 1 Introduction
$\frac{1}{2}\Big\lVert\sigma-{\mathcal{C}}_{N}\circ\cdots\circ{\mathcal{C}}_{2}\circ{\mathcal{C}}_{1}(\eta_{1}\otimes\cdots\otimes\eta_{\ell})\Big\rVert_{1}\leq\epsilon.$ (1.2)
Here $\eta_{j}$ is the initial density matrix at qudit $j$ and each ${\mathcal{C}}_{i}$ is chosen from a set of allowed elementary operations. We allow ancillary qudits inserted and discarded in between the local channels. These ancillas are all qudits of the same dimension as those in the system. In a robust definition of complexity, we should allow a constant error like $\epsilon=0.01$.
@@
In computational complexity, there is no universal definition for “elementary operations.” This ambiguity continues to exist in our study: we choose to allow elementary quantum channels to use ancillas, whereas other considerations may not; we regard arbitrary one- and two-qudit operations as elementary, while in some other cases some specific choices are preferred; we allow a constant error measured in trace distance in the approximation, while others may require smaller (more stringent) error.
@@ 2 Complexity growth of a larger region
Let $\sigma$ be a density operator obtained by tracing out a subsystem of $L$-$\ell$ qudits from a pure state on $L$ qudits. So, $\sigma$ is supported on $\ell$ qudits but the rank of $\sigma$ is bounded from above by $q^{L-\ell}$. We consider the probability that the random circuit produces two states $\sigma$ there are close to each other in fidelity. The case where $\ell=L$ was considered in [13]. Here, the near-orthogonality means that the fidelity between any pair is at most $0.1$. Our probability estimate will use a two-step argument.
@@
$\Pr_{\sigma}\big[F(\rho,\sigma)\geq 1-\delta\big]\leq\frac{\operatorname{{\mathbb{E}}}_{\sigma}F(\rho,\sigma)^{2k}}{(1-\delta)^{2k}}\leq\frac{(1+\epsilon)}{(1-\delta)^{2k}q^{k(2\ell-L)}}\min(e^{k^{2}q^{-(L-\ell)}},\penalty 10000\ k!)\,.$ (2.2)
This second step does not use any detail of the dynamics but only (2.1). Setting $k=T/\operatorname{poly}(L)$ and $\delta=0.9$, and assuming $\ell>L/2$, we conclude that the probability of interest is at most $q^{-T/\operatorname{poly}(L)}$ as long as $T\leq\min(\operatorname{poly}(L)\max(q^{2\ell-L},q^{(L-\ell)/2}),\,q^{cL})$. Note that $\rho$ is completely arbitrary.
Now, a local quantum circuit with $G$ gates, each of which is a unitary or nonunitary quantum channel, can express only $\operatorname{poly}(q,L)^{G}$ nearly orthogonal states. Therefore, the probability that the reduced density matrix $\sigma$ on a subsystem of size $\ell$ has a large fidelity to any of those $\operatorname{poly}(q,L)^{G}$ states is at most $\operatorname{poly}(q,L)^{G}q^{-T/\operatorname{poly}(L)}$. It follows that most reduced density matrices $\sigma$ produced by the random circuit cannot be expressed by $G$ gates unless
$G\geq T/\operatorname{poly}(L)\quad\text{ if }\quad T\leq e^{c^{\prime}L}$ (2.3)
for some universal constant $c^{\prime}>0$. This shows the linear-$T$ growth of the complexity of most instances $\sigma$.
In short, the argument is that the random circuits produce $N\geq q^{T/\operatorname{poly}(L)}$ nearly orthogonal states on the subsystem of size $\ell$ so we need $\log(N)\geq T/\operatorname{poly}(L)$ gates to express them. Note that the result here, especially (2.2), is meaningful only for any subsystem of size $\ell$ such that
@@
The factor of $2$ in the condition $\ell>L/2$ appears in the proof at a seemingly minor step where we bound the rank of $\sigma$ by $q^{L-\ell}$. Indeed, if the rank of $\sigma$ is much smaller than this, still the argument goes through and the linear growth of subsystem complexity holds. This leads to very different phenomenology of complexity for other random circuits, for example, the “shallow” unitary designs [33, 39, 17], which we explain in Appendix˜C.
Proof
The first inequality in (2.2) is Markov’s inequality. The second inequality is proved as follows. The fidelity between $\rho$ and $\sigma$ has its $k$-th moment bounded as
@@
$\leq r^{k}(1+\epsilon)\operatorname{{\mathbb{E}}}_{U}\operatorname{{\mathbb{E}}}_{i}(\operatorname{{Tr}}UP_{i}U^{\dagger}Q)^{k}$ $(\sigma^{\prime}=\operatorname{{\mathbb{E}}}_{i}P_{i},\penalty 10000\ \operatorname{{rank}}P_{i}=r^{\prime})$
@@
where the second to the last inequality uses the convexity of $\phi:y\mapsto y^{k}$, and the last inequality uses Lemma˜2.2 below. We finally use $r\leq q^{L-\ell}=r^{\prime}$ to arrive at the second inequality of (2.2).
@@
This is a minor generalization of [23, Lemma III.5], in which $x\mapsto\exp(\xi x)$ was used as a convex function with an optimized $\xi\in{\mathbb{R}}$. A self-contained short proof of Lemma˜2.2 is presented in Section˜A.1.
@@ 3 Late-time behavior of a smaller interval
Let $P(T;\ell)$ be the ensemble-averaged purity of a finite interval of length $\ell$ at time $T$ under the brickwork random circuit dynamics starting with a product state on an infinite chain. For any $q\geq 1$, integer $T\geq 0$, and even integer $\ell\geq 2$, we have
$\frac{1-q^{-2}}{T+1}\binom{2T}{T}\left(\frac{q}{q^{2}+1}\right)^{2T}\leq\left(P(T;\ell)-\frac{1}{q^{\ell}}\right)\leq\left(\frac{2q}{q^{2}+1}\right)^{2T}\,.$ (3.1)
@@
$\left\lVert\rho-\frac{{\mathbf{1}}_{q^{\ell}}}{q^{\ell}}\right\rVert_{1}\leq q^{\ell/2}\left\lVert\rho-\frac{{\mathbf{1}}_{q^{\ell}}}{q^{\ell}}\right\rVert_{2}=\sqrt{q^{\ell}P(T;\ell)-1}<\epsilon$ (3.2)
@@
$T>\frac{\ell}{2}\left(1+3\frac{\log 2}{\log q}\right)+\frac{3\log\tfrac{1}{\epsilon}}{\log q}\,.$ (3.3)
@@ 4.1 Mutual information growth
Let $\sigma$ be a density matrix of interest, and suppose that a circuit of local quantum channels generates a bipartite state $\rho$ such that $\frac{1}{2}\lVert\rho-\sigma\rVert_{1}=\epsilon$ where there are $m$ local channels in the circuit that straddle between complementary subsystems $A$ and $B$. Using Stinespring, we regard that all local channels are implemented by some local isometry using ancillas that are going to be traced out. The isometry in a dilation of a channel $\mathcal{C}$} can always be chosen to be a matrix product operator of bond dimension that is the product of Hilbert space dimensions of the input and output. This gives an upper bound $q^{2m}$ on the bond dimension across $A$ and $B$ for $\rho$ as a tensor network. In turn, the bond dimension bounds the von Neumann mutual information from above by $4m\log q$. Since tracing out does not increase the mutual information (4.2), we have
@@
To account for state generation error $\epsilon=\frac{1}{2}\lVert\rho-\sigma\rVert_{1}$, we use the continuity of mutual information. The inequalities of [1, 5] imply that
$|I_{\rho}(A\!:\!B)-I_{\sigma}(A\!:\!B)|\leq 10\,\epsilon\,\log(\min(d_{A},d_{B})/\epsilon)\quad\text{ assuming }\quad\epsilon<\tfrac{1}{2}$ (4.4)
@@
$I_{\sigma}(A\!:\!B)\leq 4m\log q+5\epsilon\ell\log(q/\epsilon)\,.$ (4.5)
@@ 4.2 Spatial profile
$2T\log q\geq\overline{S(\ell)}\geq\overline{S_{2}(\ell)}\geq-2T\log P(T;\ell)=2T\log\frac{q^{2}+1}{2q}>2T(1-\delta)\log q\,$, (4.7)
@@ 4.3 A complexity lower bound
$(\ell-2T)T\leq 5\epsilon\ell^{2}+4\sum_{x}m(x)\leq 5\epsilon\ell^{2}+4\mathfrak{C}(\sigma)\,.$ (4.13)
@@ 5 Complexity growth according to holographic proposals
What does holography predict for the quantum complexity of the boundary subregion as a function of time? Complexity was introduced in the context of holography in [41, 40, 10, 9]. Generalizations to the complexity of subsystems were proposed in [4, 15, 2], and applied to the time-evolved quantum quench in [14]. Roughly, different proposals for holographic complexity amount to different reasonable measures of the “size” of the entanglement wedge. Here, the entanglement wedge is a spacetime region between the HRT surface and the asymptotic boundary. Its intersection with the spatial slice is shown by the shaded regions in (5.3), and for the sake of an easy illustration, we will define the size of the entanglement wedge as the volume of this shaded region – this corresponds roughly to the “CV” prescription [4]. If we further define the complexity by subtracting the value in the thermal state, then the complexity is approximately just the volume of the gray shaded region that lies behind the horizon (dashed lines in figure (5.3)). Explicitly, for $\ell<L/2$
@@ 7.2 Using full unitary group
If we drop the condition that $U\sim\nu$ is a unitary $k$-design, and consider the full set of $\tau_{U}$ with arbitrary $U$ [21, Packing net II], then the maximum number $N_{0}$ of pairwise $\beta$-separated states $\tau_{U_{i}}$ is lower bounded by Lemma˜A.1 as, when $\beta<\alpha$,
@@
Let $P$ be a projector of rank $r$ and dimension $d$. It is not difficult to show that $\lVert UPU^{\dagger}-P\rVert_{1}\geq 2\operatorname{{Tr}}(UPU^{\dagger}({\mathbf{1}}-P))$ for any $U\in{\mathsf{U}}(d)$. Using Lemma˜A.1, we find that if $r<\epsilon d$,
@@
Note the following slightly counterintuitive aspect of this result. A quantum channel that outputs a maximally mixed state for all inputs has zero capacity. Moreover, if the output has maximal entropy minus a few bits, then the capacity is at most a few bits. In this sense, such density matrices contain little quantum information. However, the results above imply that one can find a very large number of nearly orthogonal density matrices with entropy just a few bits less than maximal. In other words, such density matrices can accurately encode the geometry of high dimensional subspaces [44, 28, 26].
@@ A.1 Projector overlap
Let $Z=\frac{d}{pq}\operatorname{{Tr}}(UPU^{\dagger}Q)$. Markov’s inequality implies that $\Pr[Z\geq 1+z]=\Pr[e^{\xi Z}\geq e^{\xi(1+z)}]\leq\operatorname{{\mathbb{E}}}_{U}e^{\xi Z}e^{-\xi(1+z)}$ and $\Pr[Z\leq 1+z]=\Pr[e^{-\xi Z}\geq e^{-\xi(1+z)}]\leq\operatorname{{\mathbb{E}}}_{U}e^{-\xi Z}e^{\xi(1+z)}$ for any $\xi>0$. Use $\phi:y\mapsto\exp(\pm\xi y)$ in Lemma˜2.2 with $\xi=pq|z|/(1+z)$, and calculate Gaussian integrals. ∎

### v2

@@ Abstract
For chaotic quantum dynamics modeled by random unitary circuits, we study the complexity of reduced density matrices of subsystems as a function of evolution time where the initial global state is a product pure state. The state complexity is defined as the minimum number of local quantum channels to generate a given state from a product state to a good approximation. In 1+11+1d, we prove that the complexity of subsystems of length $\ell$ smaller than half grows linearly in time $T$ at least up to $T=\ell/4$ but becomes zero after time $T=\ell/2$ in the limit of a large local dimension, while the complexity of the complementary subsystem of length larger than half grows linearly in time up to exponentially late times. Using holographic correspondence, we give some evidence that the state complexity of the smaller subsystem should actually grow linearly up to time $T=\ell/2$ and then abruptly decay to zero.
@@ 1 Introduction
$\frac{1}{2}\Big\lVert\sigma-{\mathcal{C}}_{N}\circ\cdots\circ{\mathcal{C}}_{2}\circ{\mathcal{C}}_{1}(\eta_{1}\otimes\cdots\otimes\eta_{\ell})\Big\rVert_{1}\leq\alpha.$ (1.2)
Here $\eta_{j}$ is the initial density matrix at qudit $j$ and each ${\mathcal{C}}_{i}$ is chosen from a set of allowed elementary operations. We allow ancillary qudits inserted and discarded in between the local channels. These ancillas are all qudits of the same dimension as those in the system. In a robust definition of complexity, we should allow a constant error like $\alpha=0.01$.
@@
In computational complexity, there is no universal definition for “elementary operations.” This ambiguity continues to exist in our study: we choose to allow elementary quantum channels to use ancillas, whereas other considerations may not; we take arbitrary two-qudit operations as elementary, while in some other cases some specific choices are preferred; we allow a constant error measured in trace distance in the approximation, while others may require smaller (more stringent) error.
@@ 2 Complexity growth of a larger region
Let $\sigma$ be a density operator obtained by tracing out a subsystem of $L$-$\ell$ qudits from a pure state on $L$ qudits. So, $\sigma$ is supported on $\ell$ qudits but the rank of $\sigma$ is bounded from above by $q^{L-\ell}<q^{\ell}$. We consider the probability that the random circuit produces two states $\sigma$ there are close to each other in fidelity. The case where $\ell=L$ was considered in [13]. Here, the near-orthogonality means that the fidelity $F$ between any pair is at most $0.1$. Our probability estimate will use a two-step argument.
@@
$\Pr_{\sigma}\big[F(\rho,\sigma)\geq 1-2\alpha\big]\leq\frac{\operatorname{{\mathbb{E}}}_{\sigma}F(\rho,\sigma)^{2k}}{(1-2\alpha)^{2k}}\leq\frac{(1+\epsilon)}{(1-2\alpha)^{2k}q^{k(2\ell-L)}}\min(e^{k^{2}q^{-(L-\ell)}},\penalty 10000\ k!)\,.$ (2.2)
This second step does not use any detail of the dynamics but only (2.1). Setting $k=T/\operatorname{poly}(L)$ and $\alpha=0.9$, and assuming $\ell>L/2$, we conclude that the probability of interest is at most $q^{-T/\operatorname{poly}(L)}$ as long as $T\leq\min(\operatorname{poly}(L)\max(q^{2\ell-L},q^{(L-\ell)/2}),\,q^{cL})$. Note that $\rho$ is completely arbitrary.
Now, let $\mathcal{S}$ be the set of all states that are expressible by $G$ gates. Suppose we have a set $\{\rho_{1},\rho_{2},\ldots,\rho_{N}\}\subset\mathcal{S}$ such that for every $\rho\in\mathcal{S}$ there exists $\rho_{i}$ with $\frac{1}{2}\lVert\rho-\rho_{i}\rVert_{1}\leq\alpha$. Such a set is called an $\alpha$-covering net for $\mathcal{S}$. If $\sigma$ is $\alpha$-close in trace distance to some element of $\mathcal{S}$, then $\sigma$ is $2\alpha$-close to some $\rho_{i}$. Hence, the probability that our reduced density matrix $\sigma$ generated by the random circuit has complexity $G$ is at most $Nq^{-T/\operatorname{poly}(L)}$. We want this probability to be so small that the complexity of $\sigma$ is typically large. Our interest is thus an upper bound on $N$.
If we perturb each gate in a circuit of $G$ gates by $\frac{\alpha}{G}$ in the channel metric (diamond norm), then the final state will differ in trace distance by at most $\alpha$. This means that for the construction of $\{\rho_{1},\ldots,\rho_{N}\}$, it is extravagant to consider two gates at a location that differ by $\frac{\alpha}{2G}$ or smaller in the channel metric. It is not difficult to construct a $\frac{\alpha}{G}$-covering net of cardinality $(G/\alpha)^{\operatorname{poly}(q)}$ for the space of all two-qudit channels, where the exponent $\operatorname{poly}(q)$ is the dimension. [^4] This covering net for the space of local channels is not needed if we declare that only $\operatorname{poly}(q)$ choices of two-qudit channels are elementary, in which case $N\leq\operatorname{poly}(q,L)^{G}$. We conclude that $N\leq\operatorname{poly}(q,L,(G/\alpha)^{\operatorname{poly}(q)})^{G}$.
It follows that most reduced density matrices $\sigma$ produced by the random circuit requires $G$ local channels where
$G\geq T/\operatorname{poly}(q,L,\log T)\geq T/\operatorname{poly}(qL)\quad\text{ if }\quad T\leq q^{c^{\prime}L}$ (2.3)
for some universal constant $c^{\prime}>0$, where the second inequality is because $\log T\leq c^{\prime}L\log q$. This shows the linear-$T$ growth of the complexity of most instances $\sigma$ until exponentially late time.
In short, the argument is that the random circuits produce $N\geq q^{T/\operatorname{poly}(qL)}$ nearly orthogonal states on the subsystem of size $\ell$ so we need $\log(N)\geq T/\operatorname{poly}(qL)$ gates to express them. Note that the result here, especially (2.2), is meaningful only for any subsystem of size $\ell$ such that
@@
The factor of $2$ in the condition (2.4) appears in the proof at a seemingly minor step where we bound the rank of $\sigma$ by $q^{L-\ell}$. The condition $\ell>L/2$ is just to ensure that $q^{L-\ell}<q^{\ell}$. Indeed, if the rank of $\sigma$ is much smaller than $q^{\ell}$ (the maximum possible), still the argument goes through and the linear growth of subsystem complexity holds. This leads to very different phenomenology of complexity for other random circuits, for example, the “shallow” unitary designs [33, 39, 17], which we explain in Appendix˜C.
Proof of Eq.˜2.2
The first inequality in Eq.˜2.2 is Markov’s inequality. The second inequality is proved as follows. The fidelity between $\rho$ and $\sigma$ has its $k$-th moment bounded as
@@
[…]
$\leq r^{k}(1+\epsilon)\operatorname{{\mathbb{E}}}_{V}\operatorname{{\mathbb{E}}}_{i}(\operatorname{{Tr}}VP_{i}V^{\dagger}Q)^{k}$ $(\rho^{\prime}=\operatorname{{\mathbb{E}}}_{i}P_{i},\penalty 10000\ \operatorname{{rank}}P_{i}=r^{\prime})$
@@
where the second to the last inequality uses the convexity of $\phi:y\mapsto y^{k}$, and the last inequality uses Lemma˜2.8 below. We finally use $r\leq q^{L-\ell}=r^{\prime}$ to arrive at the second inequality of Eq.˜2.2.
@@
This is a minor generalization of [23, Lemma III.5], in which $x\mapsto\exp(\xi x)$ was used as a convex function with an optimized $\xi\in{\mathbb{R}}$. A self-contained short proof of Lemma˜2.8 is presented in Section˜A.1.
@@ 3 Late-time behavior of a smaller interval
Let $P(T;\ell)$ be the ensemble-averaged purity of a finite interval of length $\ell$ at time $T$ under the brickwork random circuit dynamics starting with a product pure state on an infinite chain. For any $q\geq 1$, integer $T\geq 0$, and even integer $\ell\geq 2$, we have
$\frac{1-q^{-\ell}}{T+1}\binom{2T}{T}\left(\frac{q}{q^{2}+1}\right)^{2T}\leq\left(P(T;\ell)-\frac{1}{q^{\ell}}\right)\leq\left(\frac{2q}{q^{2}+1}\right)^{2T}\,.$ (3.2)
@@
$\left\lVert\rho-\frac{{\mathbf{1}}_{q^{\ell}}}{q^{\ell}}\right\rVert_{1}\leq q^{\ell/2}\left\lVert\rho-\frac{{\mathbf{1}}_{q^{\ell}}}{q^{\ell}}\right\rVert_{2}=\sqrt{q^{\ell}P(T;\ell)-1}<\alpha$ (3.3)
@@
$T>\frac{\ell}{2}\left(1+3\frac{\log 2}{\log q}\right)+\frac{3\log\tfrac{1}{\alpha}}{\log q}\,.$ (3.4)
@@ 4.1 Mutual information growth
Let $\sigma$ be a density matrix of interest, and suppose that a circuit of local quantum channels generates a bipartite state $\rho$ such that $\frac{1}{2}\lVert\rho-\sigma\rVert_{1}=\alpha$ where there are $m$ local channels in the circuit that straddle between complementary subsystems $A$ and $B$. Using Stinespring, we regard that all local channels are implemented by some local isometry using ancillas that are going to be traced out. The isometry in a dilation of a channel $\mathcal{C}$} can always be chosen to be a matrix product operator of bond dimension that is the product of Hilbert space dimensions of the input and output. This gives an upper bound $q^{2m}$ on the bond dimension across $A$ and $B$ for $\rho$ as a tensor network. In turn, the bond dimension bounds the von Neumann mutual information from above by $4m\log q$. Since tracing out does not increase the mutual information (4.2), we have
@@
To account for state generation error $\alpha=\frac{1}{2}\lVert\rho-\sigma\rVert_{1}$, we use the continuity of mutual information. The inequalities of [1, 5] imply that
$|I_{\rho}(A\!:\!B)-I_{\sigma}(A\!:\!B)|\leq 10\,\alpha\,\log(\min(d_{A},d_{B})/\alpha)\quad\text{ assuming }\quad\alpha<\tfrac{1}{2}$ (4.4)
@@
$I_{\sigma}(A\!:\!B)\leq 4m\log q+5\alpha\ell\log(q/\alpha)\,.$ (4.5)
@@ 4.2 Spatial profile
$2T\log q\geq\overline{S(\ell)}\geq\overline{S_{2}(\ell)}\geq-\log P(T;\ell)=2T\log\frac{q^{2}+1}{2q}>2T(1-\delta)\log q\,$, (4.7)
@@ 4.3 A complexity lower bound
$(\ell-2T)T\leq 5\alpha\ell^{2}+4\sum_{x}m(x)\leq 5\alpha\ell^{2}+4\mathfrak{C}(\sigma)\,.$ (4.13)
@@ 5 Complexity growth according to holographic proposals
What does holography predict for the quantum complexity of the boundary subregion as a function of time? Complexity was introduced in the context of holography in [41, 40, 10, 9]. Generalizations to the complexity of subsystems were proposed in [4, 15, 2], and applied to the time-evolved quantum quench in [14]. Roughly, different proposals for holographic complexity amount to different reasonable measures of the “size” of the entanglement wedge. Here, the entanglement wedge is a spacetime region between the HRT surface and the asymptotic boundary. Its intersection with the spatial slice is shown by the shaded regions in (5.1-5.3), and for the sake of an easy illustration, we will define the size of the entanglement wedge as the volume of this shaded region – this corresponds roughly to the “CV” prescription [4]. If we further define the complexity by subtracting the value in the thermal state, then the complexity is approximately just the volume of the gray shaded region that lies behind the horizon (dashed lines in figure (5.1-5.3)). Explicitly, for $\ell<L/2$
@@ 7.2 Using full unitary group
If we drop the condition that $U\sim\nu$ is a unitary $k$-design, and consider the full set of $\tau_{U}$ with arbitrary $U$ [21, Packing net II], then the maximum number $N_{0}$ of pairwise $\beta$-separated states $\tau_{U_{i}}$ is lower bounded by Lemma˜A.3 as, when $\beta<\alpha$,
@@
Let $P$ be a projector of rank $r$ and dimension $d$. It is not difficult to show that $\lVert UPU^{\dagger}-P\rVert_{1}\geq 2\operatorname{{Tr}}(UPU^{\dagger}({\mathbf{1}}-P))$ for any $U\in{\mathsf{U}}(d)$. Using Lemma˜A.3, we find that if $r<\epsilon d$,
@@
Note the following slightly counterintuitive aspect of this result. A quantum channel that outputs a maximally mixed state for all inputs has zero capacity. Moreover, if the output has maximal entropy minus a few bits, then the quantum capacity is at most a few bits. In this sense, such density matrices contain little quantum information. However, the results above imply that one can find a very large number of nearly orthogonal density matrices with entropy just a few bits less than maximal. In other words, channels with such outputs can have a high capacity for the classical and quantum identification problems, accurately (but not exactly) encoding the geometry of high dimensional subspaces [44, 28, 26].
@@ A.1 Projector overlap
Let $Z=\frac{d}{pq}\operatorname{{Tr}}(UPU^{\dagger}Q)$. Markov’s inequality implies that $\Pr[Z\geq 1+z]=\Pr[e^{\xi Z}\geq e^{\xi(1+z)}]\leq\operatorname{{\mathbb{E}}}_{U}e^{\xi Z}e^{-\xi(1+z)}$ and $\Pr[Z\leq 1+z]=\Pr[e^{-\xi Z}\geq e^{-\xi(1+z)}]\leq\operatorname{{\mathbb{E}}}_{U}e^{-\xi Z}e^{\xi(1+z)}$ for any $\xi>0$. Use $\phi:y\mapsto\exp(\pm\xi y)$ in Lemma˜2.8 with $\xi=pq|z|/(1+z)$, and calculate Gaussian integrals. ∎